OpenStudy (anonymous):

Which value of x makes the equation true? x – 5.8 = –4.2 x =

OpenStudy (shalayla4):

-1.6

OpenStudy (gabbyalicorn):

@shalayla4 please do not give out direct answers x−5.8=−4.2 Add 5.8 to both sides. x−5.8+5.8=−4.2+5.8 x=1.6 Answer: x=1.6

OpenStudy (anonymous):

thank both of u